# Motion-Blender Gaussian Splatting for Dynamic Scene Reconstruction

### v2

@@ Abstract
Gaussian splatting has emerged as a powerful tool for high-fidelity reconstruction of dynamic scenes. However, existing methods primarily rely on implicit motion representations, such as encoding motions into neural networks or per-Gaussian parameters, which makes it difficult to further manipulate the reconstructed motions. This lack of explicit controllability limits existing methods to replaying recorded motions only, which hinders a wider application in robotics. To address this, we propose Motion Blender Gaussian Splatting (MBGS), a novel framework that uses motion graphs as an explicit and sparse motion representation. The motion of a graph’s links is propagated to individual Gaussians via dual quaternion skinning, with learnable weight painting functions that determine the influence of each link. The motion graphs and 3D Gaussians are jointly optimized from input videos via differentiable rendering. Experiments show that MBGS achieves state-of-the-art performance on the highly challenging iPhone dataset while being competitive on HyperNeRF. We demonstrate the application potential of our method in animating novel object poses, synthesizing real robot demonstrations, and predicting robot actions through visual planning. The source code, models, video demonstrations can be found at mlzxy.github.io/motion-blender-gs.
@@ 1 Introduction
Reconstructing and modeling dynamic 3D scenes is a fundamental challenge in computer vision, with applications on content creation, augmented reality, and robotics. Recent work on 3D Gaussian splatting has made significant progress in capturing dynamic scenes, enabling efficient and high-fidelity reconstruction [1, 2, 3, 4, 5, 6, 7]. Existing Gaussian splatting methods utilize 3D Gaussians to represent geometry and appearance, paired with motion modules that determine the movements of the Gaussians. For instance, 4D-Gaussians [3] and Deformable-GS [8] encode motion implicitly into neural networks. Shape-of-Motion [1] and STG [4] use shallower models that require dense per-Gaussian motion parameters. While these approaches achieve high-fidelity reconstruction, the reconstructed scenes cannot be easily altered or manipulated in simulation due to their implicit motion representation, which limits their use for robot manipulation planning. This lack of direct and explicit controllability restricts existing methods to simply replaying recorded motions.
@@ 2.1 Motion Blender Gaussian Splatting
Our framework is designed based on two key ideas that set it apart from existing works. 111 Represent motions with a sparse structure: Instead of associating dense motion parameters to each Gaussian (often numbering in the hundreds of thousands), we use a motion graph with far fewer parameters (often less than 100). The sparse link-wise motions are then propagated to each Gaussian through motion blending. 222 Represent motions explicitly: Instead of encoding motions implicitly into neural networks, our motion graph is an explicit kinematic model. This allows a straightforward visualization and manipulation of motions in the 3D space, such as transferring motion patterns from one object to a new object, or creating novel animations through motion graph editing.
@@
$\mathcal{R}:\operatorname{SE}(3)^{|\mathcal{L}|}\times\operatorname{SE}(3)^{|\mathcal{L}|}\mapsto\operatorname{SE}(3)^{|\mathcal{L}|}$ returns relative rigid transforms between two sets of 3D poses. Specifically, $\mathcal{R}\left(P_{\mathcal{L},0},P_{\mathcal{L},t}\right)$ indicates the SE(3) movement from time 00 to time $t$ of each link in $\mathcal{L}$.
@@ 3.3 Applications
Figure 10: Robotic Action Prediction via Visual Planning. We reconstruct object Gaussians (cloth, microwave, rope) from human videos, simulate potential motion graph trajectories, and render the resulting object images. The optimal trajectory is selected by maximizing PSNR between rendered images and the target goal image. Key joint trajectories (yellow) are mapped to end-effector positions. Videos of our robot experiments can be found at mlzxy.github.io/motion-blender-gs.
@@ A.1 Related Works
Explicit Motion Representations. Explicit motion representations offer intuitive visualization and straightforward motion manipulation. Gaussian-based avatar modeling methods [53, 54, 55] rely on parametric human templates, such as SMPL [56], but are inherently limited to human subjects and cannot generalize to arbitrary dynamic objects. SC-GS [7] introduces sparse control points, which can be dragged to create new motions. But SC-GS is restricted to simple, synthetic scenes like those in D-NeRF [50]. CAMM [57] employs kinematic chains for motion representation but is limited to meshes instead of 3D Gaussians and requires occlusion-free videos with clean backgrounds. On the other hand, classical animation techniques [9] utilize explicit motion representations such as deformation graphs [10], harmonic coordinates [11], and cage [12]. However, these methods only focus on applying manually designed motions to mesh surfaces and are not designed for reconstructing motion or geometry from unstructured video data. In contrast, our method introduces a graph-based motion representation for Gaussian splatting, where motion is propagated from sparse graph links to individual Gaussians. 3D Gaussians and motion graphs are jointly optimized from videos of realistic and complex scenes.
@@ A.3 More Details of Real Robot Experiments
For the green cloth, we initialize the motion graph at $t=0$ using a rectangular mesh generated via NanoMesh [67] (Fig. 10, first column). For the microwave and rope—which lack standardized keypoint structures—we manually define their kinematic trees. The microwave is represented as a 2-link graph, with 3D keypoints annotated in the first ($t=0$) and final ($t=T-1$) frames. For the rope, we annotate ten 2D keypoints at the first frame, and track them across frames using CoTracker [68]. The rope’s kinematic tree is initialized by projecting the ten manually annotated 2D keypoints at the first frame into 3D space using depth information, with the central keypoint serving as the root. During optimization, we apply 2D keypoint regularization to the rope but omit it for the cloth and microwave.

### v4

@@ Abstract
Gaussian splatting has emerged as a powerful tool for high-fidelity reconstruction of dynamic scenes. However, existing methods primarily rely on implicit motion representations, such as encoding motions into neural networks or per-Gaussian parameters, which makes it difficult to further manipulate the reconstructed motions. This lack of explicit controllability limits existing methods to replaying recorded motions only, which hinders a wider application in robotics. To address this, we propose Motion Blender Gaussian Splatting (MBGS), a novel framework that uses motion graphs as an explicit and sparse motion representation. The motion of a graph’s links is propagated to individual Gaussians via dual quaternion skinning, with learnable weight painting functions that determine the influence of each link. The motion graphs and 3D Gaussians are jointly optimized from input videos via differentiable rendering. Experiments show that MBGS achieves state-of-the-art performance on the highly challenging iPhone dataset while being competitive on HyperNeRF. We demonstrate the application potential of our method in animating novel object poses, synthesizing real robot demonstrations, and predicting robot actions through visual planning. The source code and the models are included in the supplementary material. Video demonstrations can be found at mlzxy.github.io/motion-blender-gs.
@@ 1 Introduction
Reconstructing and modeling dynamic 3D scenes is a fundamental challenge in robot vision. Recent work on 3D Gaussian splatting has made significant progress in capturing dynamic scenes, enabling efficient and high-fidelity reconstruction [1, 2, 3, 4, 5, 6, 7]. Existing Gaussian splatting methods utilize 3D Gaussians to represent geometry and appearance, paired with motion modules that determine the movements of the Gaussians. For instance, 4D-Gaussians [3] and Deformable-GS [8] encode motion implicitly into neural networks. Shape-of-Motion [1] and STG [4] use shallower models that require dense per-Gaussian motion parameters. While these approaches achieve high-fidelity reconstruction, the reconstructed scenes cannot be easily altered or manipulated in simulation due to their implicit motion representation, which limits their use for robot manipulation planning. This lack of direct and explicit controllability restricts existing methods to simply replaying recorded motions.
@@ 2.1 Motion Blender Gaussian Splatting
Our framework is designed based on two key ideas that set it apart from existing works. 1 Represent motions with a sparse structure: Instead of associating dense motion parameters to each Gaussian (often numbering in the hundreds of thousands), we use a motion graph with far fewer parameters (often less than 100). The sparse link-wise motions are then propagated to each Gaussian through motion blending. 2 Represent motions explicitly: Instead of encoding motions implicitly into neural networks, our motion graph is an explicit kinematic model. This allows a straightforward visualization and manipulation of motions in the 3D space, such as transferring motion patterns from one object to a new object, or creating novel animations through motion graph editing.
@@
$\mathcal{R}:\operatorname{SE}(3)^{|\mathcal{L}|}\times\operatorname{SE}(3)^{|\mathcal{L}|}\mapsto\operatorname{SE}(3)^{|\mathcal{L}|}$ returns relative rigid transforms between two sets of 3D poses. Specifically, $\mathcal{R}\left(P_{\mathcal{L},0},P_{\mathcal{L},t}\right)$ indicates the SE(3) movement from time 0 to time $t$ of each link in $\mathcal{L}$.
@@ 3.2 Results
Figure 7: Novel view rendering on iPhone with best LPIPS in red. More results in Fig. A4.
@@
Figure 8: Visualization on HyperNerf with best LPIPS in red. More results in Fig. A5
@@ 3.3 Applications
Figure 10: Robotic Action Prediction via Visual Planning. We reconstruct object Gaussians (cloth, microwave, rope) from human videos, simulate potential motion graph trajectories, and render the resulting object images. The optimal trajectory is selected by maximizing PSNR between rendered images and the target goal image. Key joint trajectories (yellow) are mapped to end-effector positions. Videos of our robot experiments can be found at anonymous368.github.io/motion-blender.
@@ 4 Final Remarks
5 Limitations and Future Directions
Visual Artifacts on Novel Poses. Fig. 11 (b) demonstrates failure cases in animating novel object poses, where attempting to turn the cat’s head and rotate the windmill’s badge introduces non-negligible visual artifacts. Visual artifacts can also be found in Fig. 9 and Fig. 10 of our robot experiments. We believe the key reason is that, unlike meshes, Gaussians lack an explicit surface representation. This allows each individual Gaussian to deform arbitrarily when motion graphs are applied. This can cause some Gaussians to deviate from object surfaces, particularly for unseen motions. Recent advances like 2D-GS [38], which directly formulates Gaussians on 3D surfaces, offer promising solutions to this limitation.
Figure 11: Failure Cases. Visualization of imperfect learned motion graphs (a), failure cases of novel pose editing (b), and failure cases of reconstructing reflective surfaces (c).
Ungrounded Motion Graphs. Fig. 11 (a) demonstrates imperfect learned motion graphs that are not accurately grounded to the object geometry. Top left: the hand contains intricate kinematic structures but only occupies a small region (less than $80\times 80$ pixels in a $720\times 920$ image), making it difficult to capture fine details. Top right: incomplete motion graph coverage on the cat’s right front leg and joints protruding beyond the cat body. Bottom left: the pillow is squeezed but the motion graph fails to sufficiently deform. Bottom right: the kinematic tree does not accurately align with the microwave door.
We believe the key reason is that Gaussian splatting reconstructs scenes purely from visual observations, without incorporating structural or physical priors. For instance, Fig. 12 illustrates a failure case in rope manipulation. While the reconstructed motion graph appears accurate in 2D image space, a 3D inspection reveals that the rope’s head tilts below the table—an incorrect reconstruction undetectable from visual observation alone. Since the robot’s end-effector pose is derived from the motion graph’s joints and links, this leads to an erroneous orientation in the planned trajectory. However, recovering sparse structures from only a few videos is a highly under-constrained optimization problem. A promising direction is to incorporate semantic or physical priors from foundation models to improve geometric consistency.
Figure 12: A failure case of incorrect gripper orientation caused by an erroneous motion graph. The head of the rope motion graph is misaligned, tilting downward below the table surface.
Fast Camera and Object Motions. Fig. 13 shows failure cases when reconstructing fast-moving objects. The left figure compares our MB-GS with Shape-of-Motion on adjacent frames from the HyperNerf peel-banana scene, which is captured with strong camera motions and objects (the hand and banana) frequently entering and exiting the view. In this setting, MB-GS produces inconsistent results across frames such as a shaking hand. The right figure compares MB-GS reconstructions of a microwave with doors moving at different speeds. We observe that reconstruction quality degrades significantly when the door moves rapidly. Dynamic reconstruction under large motions remains a longstanding challenge [3]. Shape-of-Motion attempts to address this by using 2D tracks [39] as motion guidance. However, 2D tracks are often unreliable, especially under orientation changes, such as when the microwave doors rotate and original tracks on the door surface are lost. Another limitation is that MB-GS learns motion graph parameters independently at each frame, making it less effective at leveraging temporal continuity across frames. A promising direction is to still follow the motion blender framework but introduce neural networks that predict the deformation of motion graphs over time, preserving the manipulability of the representation while improving learning capacity.
Figure 13: Failure Cases on Fast-Moving Objects. Visualization of shaking artifacts from strong camera motion (left), and reconstruction quality comparison of our MB-GS for the same microwave with doors moving at different speeds (right).
Robotic Applications. In the following, we highlight two promising directions that are specific to the robotic applications of our method MB-GS.
  • Motion Graphs and Physics Simulation. Our framework synthesizes robotic demonstrations for tasks like grasping toys and folding clothes (Fig. 9), and performs planning in graph space to reach goal configurations (Fig. 10). However, our current motion graph representation lacks physical awareness. This is evident in Fig. 9 (second row), where the gripper penetrates the cloth in a physically unrealistic manner. Additionally, the graph-space planning is limited to simple shape (2D rectangle) and basic kinematic chain (chain-like graph), whose motion patterns can be reasonably approximated with geometric heuristics.
    Our sparse motion graph design naturally extends to physics-based simulation frameworks. For instance, MuJoCo [40] models deformable objects using flex elements (triangles in 2D, tetrahedra in 3D), which aligns closely with our deformable graph representation. To further bridge the gap, it is also possible to, first, incorporate not only link poses but also triangular face poses to guide Gaussian motion, leveraging projection points analogous to Fig. 3; second, support hybrid deformable and kinematic structures through a more systematic engineering integration. These extensions would significantly improve simulation compatibility for both articulated and deformable objects.
    In doing so, instead of learning joint positions or rotation angles as in our current approach, it is possible to learn MuJoCo’s physical parameters through differentiable simulation: simulating the structure, rendering images, and back-propagating through the physics engine. This would be particularly feasible in controlled environments with clear object segmentation, enabling more powerful digital twins that handle deformable and articulated objects — a capability lacking in current rigid-body-focused digital twins [41].
  • Modeling of Light Sources and Reflective Surfaces. Robotic workspaces typically contain strong light sources, and robot arms often have reflective surfaces. However, as shown in Fig. 11, our reconstructed robot surfaces appear blurry and lose their reflectivity. This occurs because Gaussian splatting does not account for lighting effects during reconstruction. While this limitation is acceptable for vanilla Gaussian splatting in static scenes (where lighting effects remain mostly constant), it becomes problematic for dynamic scenes. In such cases, surfaces under varying lighting conditions appear to move constantly, leading to blurry reconstructions. Furthermore, workspace lighting affects not only robot reconstruction but also object reconstruction, as objects may be reflective or cast different shadows when illuminated from various angles. We find that increased lighting generally results in blurrier reconstructed objects. While this is a fundamental limitation of current Gaussian splatting methods rather than specific to our approach, it highlights the need for reflection-aware Gaussian splatting techniques to improve robotic applications.
@@ A.1 Related Works
Explicit Motion Representations. Explicit motion representations offer intuitive visualization and straightforward motion manipulation. Gaussian-based avatar modeling methods [53, 54, 55] rely on parametric human templates, such as SMPL [56], but are inherently limited to human subjects and cannot generalize to arbitrary dynamic objects. SC-GS [7] introduces sparse control points, which can be dragged to create new motions. But SC-GS is restricted to simple, synthetic scenes like those in D-NeRF [50]. MoSca [57] employs a scaffold structure to represent motions and deformations. CAMM [58] employs kinematic chains for motion representation but is limited to meshes instead of 3D Gaussians and requires occlusion-free videos with clean backgrounds. On the other hand, classical animation techniques [9] utilize explicit motion representations such as deformation graphs [10], harmonic coordinates [11], and cage [12]. However, these methods only focus on applying manually designed motions to mesh surfaces and are not designed for reconstructing motion or geometry from unstructured video data. In contrast, our method introduces a graph-based motion representation for Gaussian splatting, where motion is propagated from sparse graph links to individual Gaussians. 3D Gaussians and motion graphs are jointly optimized from videos of realistic and complex scenes.
@@ A.3 More Details of Real Robot Experiments
For the green cloth, we initialize the motion graph at $t=0$ using a rectangular mesh generated via NanoMesh [68] (Fig. 10, first column). For the microwave and rope—which lack standardized keypoint structures—we manually define their kinematic trees. The microwave is represented as a 2-link graph, with 3D keypoints annotated in the first ($t=0$) and final (t=T1t=T-1) frames. For the rope, we annotate ten 2D keypoints at the first frame, and track them across frames using CoTracker [69]. The rope’s kinematic tree is initialized by projecting the ten manually annotated 2D keypoints at the first frame into 3D space using depth information, with the central keypoint serving as the root. During optimization, we apply 2D keypoint regularization to the rope but omit it for the cloth and microwave.